# Can Large Language Models Capture Video Game Engagement?

### v1

@@ V-C Text-based Engagement Prediction
We started this section with a hypothesis that using text-descriptions of frames would improve predictive capacity of LLMs compared against multimodal inputs. We believed this would be the case because the generation of text descriptions would act as a type of game-agnostic encoding, putting more emphasis on the layout and action of frames. The results presented here indicate that this is not the case. In general, obtained results show no significant differences between the Text Input and Multimodal Input - 1 Image setups. Generating text-descriptions first and using text-only input cannot provide a better encoding than simpler multimodal approaches for this task. The main reason behind this probably lies within how these models handle vision input. While LLaVA relies on CLIP for image encoding, GPT-4o uses a custom multimodal end-to-end architecture. Because both models were trained for to encode images and text into a shared embedding space [58] the extra “image to text” step is unnecessary.
V-D Multi-Image One-Shot and Few-Shot Prompting
Figure 7: $\Delta A$ values of GPT-4o models across games and across different experimental settings. The bottom two rows depict the $\Delta A$ values of the Multimodal Input - 2 Image, One-Shot and Few-Shot experiments. The rightmost column shows $\Delta A$ values averaged across all games.
In this section we present experiments using Multimodal Input - 2 Image, One-Shot and Few-Shot strategies (see Fig. 2 and Section IV for more details on these approaches). In these experiments we opt to employ the GPT-4o model only; the reason for doing so is two-fold. First, GPT models have been observed to be more consistent and perform better across all games in experiments presented in the previous sections. Second, models of the LLaVA family are limited in how they can process images as input. As mentioned in Section IV-B LLaVA models can only take single images in their input space, while GPT-4o uses a tile-based input tokenizer that is able to handle multiple images.
Figure 7 presents the results of our Multimodal Input - 2 Image, One-Shot and Few-Shot experiments compared to the best overall Text Input and Multimodal Input - 1 Image results obtained using GPT-4o. We can see that the best overall performance is achieved when using Few-Shot prompting. While the relative improvement over the baseline is not significant across all games, there is a clear pattern of improvement compared to other models. Comparing results between Fig. 6 and Fig. 7, can see that GPT-4o models significantly outperform LLaVA models on several experimental setups. These setups include LLaVA-7b on the Multimodal Input - 1 Image (Stitched) task; LLaVA-13b models on both Text Input setups; and LLaVA-34b on the Text Input - Basic Description and Multimodal Input - 1 Image (Stitched) tasks. While there is no significant difference between one-shot and few-shot prompting, the latter strategy improves the performance in $13$ out of $20$ experimental settings. We note only $6$ out of $20$ settings where the introduction of few-shot prompting decreased the performance. Our findings are aligned with results reported in the literature [63, 66, 64] suggesting that a few-show, multimodal, chain-of-thought prompting method can significantly improve LLM performance. However even with this performance boost, the observed models barely surpass the majority baseline, on average, across games.
Looking at the best and worst performances of GPT-4o across games we observe a familiar pattern. Once again, the games whose engagement is easier to predict are Wolfram ($38\%$), Apex Legends ($27\%$), and Doom ($23\%$) when we look at the average performance across both the GPT-4o One-Shot and Few-Shot settings. Similarly, the games where the LLM models performed worst on average are Counter Strike 1.6 ($-40\%$), Corridor 7 ($-25\%$), and Heretic ($-17\%$). These findings are in line with our previous experiments.
@@ V-E Qualitative Analysis
In this section we outline the reasons for the observed poor performance of the tested LLMs and analyse why certain games are easier to predict. For our analysis we are looking at the highest performing model, the GPT-4o with Multimodal Input - 2 Images using Few-Shot prompting. Employing this model we list $5$ games where the $\Delta A$ exceeds $25\%$: Doom, Wolfram, PlayerUnknown’s Battlegrounds (PUBG) (2018), Apex Legends, and Borderlands 3 (2019). Conversely, the five games, where the performance was well-below the baseline are as follows: Corridor 7, Heretic, Counter Strike 1.6, Medal of Honor (2010), and HROT; see Fig. 8.
@@
[…]
@@ VI Discussion
The evaluation experiments presented in this paper are the first of its kind for LLM-based engagement prediction in games. While collectively we tried $2,440$ combinations of experimental settings—varying the LLM model type, model size, prompting strategy, input type, and ground truth processing—there are still many aspects that we did not explore in this initial study. We argue, however, that we set out to lay ground works for future research by approaching the problem of automating gameplay annotation in a relatively straightforward way. While, for instance, we experimented with several out-of-box LLM models and prompting strategies, we kept the granularity of the vision input constant which potentially poses a core limitation to this initial study. Since we sample the videos in question at a 3-second interval, the model loses a lot of information between these frames. Although we briefly experimented with different time intervals (i.e. between 1 and 5 seconds), simply increasing the sampling rate did not yield a performance increase. It is likely, however, that by either providing more frames per query or using video input directly would lead to a significant performance improvement that remains to be tested in future studies. These investigations were purposefully left out of the scope of the current study, mainly because (at the time of writing) there were no widely available video models which could have have fit into the experimental protocol presented here.
@@
While video input could feed more information to the LLM, the context of the query could also be augmented, and then provided to the LLM, thereby improving its predictive capacity. By implementing a memory mechanism [67], for instance, we could potentially store and recall the temporal context of the play session, providing richer information to the model. Similarly, we could provide more context on the necessary domain knowledge for the task by implementing retrieval-augmented generation [68], where we could feed more information on the game, play session, or downstream task similarly to how we have been providing positive and negative examples to the model in our few-shot examples. We plan to pursue these avenues in our future studies in our effort to further investigate how more contextual information impacts the performance of LLMs towards fully autonomous engagement annotation.
@@
Generating subjective labels is a relatively open field with a lot of unanswered questions. Naturally, the exploration should be extended into other datasets, involving games—also beyond first-person shooters—and other media as well. While this study focuses on engagement, there are other subjective aspects of both player and viewer experience that could be evaluated further. A natural step forward would be to make use of a diverse set of affective corpora, focusing, for instance, on affect prediction across videogame datasets [69], but also architectural spaces [70] and movie corpora [71]. As discussed above, the current evaluation of LLMs—even though it was multimodal—considered a predetermined number of modalities: text and images. As we move forward and more multimodal architectures become widely adopted, the research into utilizing LLMs for autonomous affect annotation could encompass different modalities from images, through video, to audio. When it comes to interactive mediums such as games, user behavioural data could also be included [31] providing a richer context to the models.
@@ VII Conclusion
This paper explored a novel application of LLMs for autonomously annotating the continuous experience of viewers when consuming videos of first-person shooter video games from the GameVibe corpus. We conducted an in-depth analysis comparing multiple foundation models, including Open-AI’s GPT and the LLaVa model families, and evaluated their performance across different input modalities (i.e. multimodal, text-based) and prompting strategies (i.e. one-shot, few-shot). Our findings confirm that model size and prompting strategy have a critical impact on model performance. The LLMs presented here demonstrate promising capabilities on certain game elicitors—although their overall performance only marginally surpasses the baseline. Perhaps unsurprisingly, the games where LLMs are successful predicting the continuos change in engagement are popular games with easy-to-read graphical styles and concise gameplay. The gap in performance on these games compared to more challenging elicitors shows that while LLMs have potential, there is still a long road ahead towards automated continuous affect labelling using using these type of foundation models.
@@
Figure 10: Multimodal Input - 1-Image (Stitched) input and output example using GPT-4o.
Figure 11: Text Input - Basic Description image-to-text description generation using GPT-4o.
Figure 12: Text Input - Advanced Description image-to-text description generation using GPT-4o.
Figure 13: Text Input - Engagement Prediction using GPT-4o. $<$description A$>$ and $<$description B$>$ are replaced with generated text descriptions.
Figure 14: Multimodal Input - 2-Image One-Shot input and output example using GPT-4o.
Figure 15: Multimodal Input - 2-Image Few-Shot Example Reasoning Generation input and output example using GPT-4o. The generated output is used as an artificial example in the Few-Shot experiments.
Figure 16: Multimodal Input - 2-Image Few-Shot Engagement Prediction input and output example using GPT-4o. The Base and Task Prompts are identical to the ones described in Figure 14. Example Outputs are generated as shown in Figure 15.

### v2

@@ V-C Engagement Prediction Based on Text Descriptions
We started this section with a hypothesis that using text-descriptions of frames would improve predictive capacity of LLMs compared to multimodal inputs. We believed this would be the case because the generation of text descriptions would act as a type of game-agnostic encoding, putting more emphasis on the layout and action of frames. The results presented here indicate that this is not the case. In general, obtained results show no significant differences between the Text Descriptions and Multimodal One-Shot setups. Generating text-descriptions first and using text-only input cannot provide a better encoding than simpler multimodal approaches for this task. Because Multimodal LLMs were trained to encode images and text into a shared embedding space [58] the extra “image to text” step is unnecessary.
V-D Multimodal Few-Shot Prompting
Figure 7: Results of the Multimodal Few-Shot experiments across all models compared to the best Multimodal One-Shot results from Section V-B. The last column show the average $\Delta A$ across all games. The last two rows show the MLP and CNN Baselines.
In this section we present experiments using Multimodal Few-Shot strategy (see Fig. 2 and Section IV for more details on this approach). As there is no significant difference between the Multimodal One-Shot experiments and those that are using Text Descriptions, here we only focus on comparisons between the Multimodal setups. Figure 7 presents the results of our Multimodal Few-Shot experiments compared to the best Multimodal One-Shot results presented in Section V-B. We can observe an overall improvement of performance in the Few-Shot setup.
Similarly to the experiments based on Text Descriptions, we make the following comparisons: 1) Few-Shot experiments against each other; 2) Few-Shot experiments against One-Shot and Baseline experiments within the same model; 3) performance on individual games across all Few-Shot experiments; and finally 4) performance on individual games between Few-Shot and One-Shot setups. In the first two cases, our degree of freedom is 19, while in the last two cases it is 5. We make 15, 8, 190, and 20 comparisons per test, respectively.
When we compare the average performance of models on the Few-Shot setup, we observe that GPT-4o outperforms all other models. However, when we compare the results across games, this outstanding performance is only significant when comparing it to InternVL2.5, $t(19)=4.4$, $p=0.015$. Comparing models between tasks (Few-Shot, One-Shot, and Baselines), the performance differences between Few-Shot and One-Shot tasks are not significant. Nevertheless, GPT-4o significantly outperforms both the MLP and CNN Baselines, $t(19)=4.3$, $p=0.005$ and $t(5)=4.2$, $p=0.005$ respectively. Interestingly, while we can see that games which have been easy or hard to predict in the One-Shot setting are consistent in the Few-Shot setting, there is no significant difference between games. Nevertheless, we can observe that once again, Wolfram, Blitz Brigade, and Doom define the easiest downstream tasks; and Counter Strike 1.6, Medal of Honor, and Corridor 7 are among the hardest to predict. Comparing task performance within single games, we can observe improvement when going from One-Shot to Few-Shot prompting. This is especially true in the otherwise difficult-to-predict games: Corridor 7 ($17\%$); HROT ($13\%$); and Counter Strike 1.6 ($12\%$). Nevertheless, the only game where the performance difference across all models is significant is CS:GO - Dust2 (2012), $t(5)=2.9$, p=0.034p=0.034—where we see a $7\%$ drop.
While there are few significant differences between the One-Shot and Few-Shot experiments, there is a clear pattern of overall improvement when it comes to GPT-4o, Gemma 3, and Qwen2.5-VL. Most notably GPT-4o achieves an average $6\%$ $\Delta A$ and Qwen2.5-VL shows $59\%$ $\Delta A$ on Wolfram—the absolute best result across all experiments. However, similarly to previous experiments, there are also negative outliers. Performance in Doom is dropped by $10\%$; and in case of LLaVA-OV we can see an overall $5\%$ drop between One-Shot and Few-Shot setups. Because there is no clear pattern between models, differences must come down to the specifics of the data and the architectures of individual models.
@@ V-E Qualitative Analysis
In this section we outline the reasons for the observed poor performance of the tested LLMs and analyse why certain games are easier to predict. For our analysis we are looking at the highest performing model, the GPT-4o with Multimodal Few-Shot prompting. Employing this model we list $5$ games where the $\Delta A$ exceeds $25\%$: Doom, Wolfram, PlayerUnknown’s Battlegrounds (PUBG) (2018), Apex Legends, and Borderlands 3 (2019). Conversely, the five games, where the performance was well-below the baseline are as follows: Corridor 7, Heretic, Counter Strike 1.6, Medal of Honor (2010), and HROT; see Fig. 8.
@@
Another way to explain the fluctuation in LLM performance is the familiarity of the model with the games per se. We observe that more popular games (such as Counter Strike, Apex Legends, and PUBG, with a peak viewership [^14] Peak viewership refers to the historically highest number of concurrent viewers watching a stream. It is indicative of the maximum audience size. of $1,914,861$, $674,070$, and $597,663$, respectively on Twitch [^15] Numbers retrieved from https://twitchtracker.com/, January 2025.) yield generally better engagement predictions compared to less popular games (such as HROT, Heretic, and Corridor 7, with a peak viewership of $24,721$, $2,280$, and $195$ on Twitch15), although we should be careful with naive over-generalizations from these findings. For one, Counter Strike 1.6 is a variant of a very popular game with a peak viewership of $125,378$ in itself, but the models struggle with correctly evaluating the change in engagement—at least in the GameVibe dataset. While it is possible that the training data of GPT-4o contains images from more popular games, attempting to verify this by reconstructing parts of the GPT-4o training data is out of the scope of this paper.
@@ VI Discussion
The evaluation experiments presented in this paper are the first of its kind for LLM-based engagement prediction in games. While collectively we tried $4,840$ combinations of experimental settings—varying the model type, prompting strategy, input type, game, and ground truth—there are still many aspects that we did not explore in this initial study. We argue, however, that we set out to lay ground work for future research by approaching the problem of automating gameplay annotation in a relatively straightforward way.
VI-A Limitations
The study presented here constitutes a preliminary look into how out-of-box LLMs can solve affective computing tasks relying on their knowledge priors. Naturally, the results of such a study face multiple limitations.
VI-A1 Static Images
@@
We experimented with several models and prompting strategies but kept the granularity of the vision input constant. We focused on single frames sampled at a 3-second interval because the low resolution input is less costly and more readily available than high frames-per-second videos. Nevertheless, there is a high likelihood that some of the context of the action is lost during this processing. This is a core limitation of this initial study. Although we briefly experimented with different time intervals (i.e. between 1 and 5 seconds), simply increasing the sampling rate did not yield a performance increase. It is likely, however, that by either providing more frames per query or using video input directly would lead to a significant performance improvement that remains to be tested in future studies.
VI-A2 No Fine-Tuning
As our focus was to evaluate the state of LLMs relying solely on their own knowledge priors, we did not employ any form of fine-tuning. The main motivation for this decision is that we are interested in the zero-shot capabilities of available LLMs and their limits using prompt-engineering. While we can expect models with specialised fine-tuning to outperform general models, this is not always the case. Large commercial LLMs using few-shot prompting have shown greater flexibility and superior performance compared to pre-trained models that have been fine-tuned on specialised datasets in certain domains [77]. As full-parameter fine-tuning is prohibitively costly, it is worthwhile investigating other avenues of making use of these models. Here we demonstrated prompt-engineering including a few-shot prompting strategy, but future studies should investigate efficient parameter tuning, such as low-rank adaptation [49, 54].
VI-A3 Limited Domain
We showcase our results on $20$ different games with varying aesthetics and play-styles, however, the experiments focus only on the first-person shooter genre. Even though this is a popular videogame genre—with countless examples out there, we show that positive results generalise poorly even within this limited domain. Our results show that clear visuals are key to the models’ success. This holds true across different experiments and architectures. These results show the need for generalised game representation beyond simple pixels on the screen [71].
VI-B Future Work
@@
While video input could feed more information to the LLM, the context of the query could also be augmented and then provided to the LLM, thereby improving its predictive capacity. By implementing a memory mechanism [89], we could potentially store and recall the temporal context of the play session. Similarly, we could provide more context on the necessary domain knowledge for the task by implementing retrieval-augmented generation [33], where we could feed more information on the game or downstream task. We plan to pursue these avenues in our future studies in our effort to further investigate how more contextual information impacts the performance of LLMs towards fully autonomous engagement annotation.
@@
Generating subjective labels is a relatively open field with a lot of unanswered questions. Naturally, the exploration should be extended to other datasets. While this study focusses on engagement, there are other subjective aspects of both player and viewer experience that could be evaluated further. A natural step forward would be to make use of a diverse set of affective corpora, focusing, for instance, on affect prediction across videogame datasets [46], but also architectural spaces [79] and movie corpora [27]. As discussed above, the current evaluation of LLMs—even though multimodal—considers a predetermined number of modalities: text and images. As we move forward and more multimodal architectures become widely adopted, the research into utilizing LLMs for autonomous affect annotation could encompass different modalities from images, through video, to audio. When it comes to interactive media such as games, user behavioural data could also be included [62] providing a richer context to the models.
@@ VII Conclusion
This paper explored a novel application of LLMs for autonomously annotating the continuous experience of viewers when consuming videos of first-person shooter video games from the GameVibe corpus. We conducted an in-depth analysis comparing multiple open-source foundation models—comparing them to Open-AI’s GPT—and evaluated their performance across different input modalities (i.e. multimodal, text-based) and prompting strategies (i.e. one-shot, few-shot). While results obtained are far from what one could consider as good performance, all LLMs examined surpass the performance of our traditional ML baseline methods across all experiments. This shows the potential of LLMs in otherwise complex affect modelling tasks. Our findings confirm that model size and prompting strategy have a critical impact on model performance. The LLMs presented here demonstrate promising capabilities on certain game elicitors—although their overall performance only marginally surpasses the baseline. Perhaps unsurprisingly, the games where LLMs are successful at predicting the continuos change in engagement are popular games with easy-to-read graphical styles and concise gameplay. The gap in performance on these games compared to more challenging elicitors shows that while LLMs have potential, there is still a long road ahead towards automated continuous affect labelling using these type of foundation models.
@@
Acknowledgment
This research has been supported by the Foundation.AI project funded by the Maltese Ministry for Education, Sport, Youth, Research and Innovation (MEYR), in consultation with the University of Malta (UM) and the Malta Chamber of Commerce, Enterprise and Industry (MCCEI).